A string breaks if it`s tension exceeds 10 Newtons. A stone of mass 250 gm tied to a string of length 10 cm is rotated in a horizontal circle. The maximum angular velocity of rotation can be-(1) 20 rad/s(2)40 rad/s(3)100 rad/s(4)200 rad/s

Grade:11

1 Answers

Vikas TU
7 years ago
Since it is rotating in a horizontal frame gravity doen’t works here.
Hence balancing the horizontal forces we get,
T = m*w^2*l
10 = (250/1000)* w^2*(10/100)
w^2 = 400
w = 20 rad/s.